OpenStudy (shamil98):

f(x) = 1 - cos x / x^2 + 1 Quotient rule. f(x) = 1 - cos x f'(x) = 1 + sin x g(x) = x^2 + 1 g'(x) = 2x 1 + sin x (x^2 + 1) - 2x(1-cosx) / (x^2 + 1 )^2 Am I correct?

OpenStudy (shamil98):

oops, made a mistake..

OpenStudy (luigi0210):

Yup

OpenStudy (shamil98):

sin x (x^2 + 1) - 2x(1-cosx) / (x^2 + 1 )^2 there is no 1 +..

OpenStudy (anonymous):

Not Quite f'(x) should just be Sin(x)

OpenStudy (anonymous):

But otherwise that is right.

OpenStudy (shamil98):

thanks.

OpenStudy (anonymous):

\[f'\left( \frac{ u }{ v } \right)=\frac{ vu'-uv' }{v ^{2} }\]

OpenStudy (anonymous):

\[f'\left( \frac{ 1-\cos x }{ x ^{2}+1 } \right)=\frac{ \sin x \left( x ^{2}+1 \right)-2x \left( 1-\cos x \right) }{\left( x ^{2} +1 \right)^{2} }\] you are absolutely correct.